Monochromatic light from a distant source is incident on a slit 0.850 mm wide. On a screen 2.11 m away, the distance from the central maximum of the diffraction pattern to the first minimum is measured to be 1.38 mm .

Calculate the wavelength of the light.
Answer #1

For the first minimum.

wy/D = λ

=> λ = (0.85 * 10-3) * (1.38 * 10-3) / 2.11 = 5.56 * 10-7 m